Write the polynomial function of minimum degree in standard form with real coefficients whose zeros include 1,1-i ,1=i ?


1
Expert's answer
2021-10-07T15:14:40-0400


"P(x)=a(x-x_1)(x-x_2)(x-x_3)"

Let "a=1." Then


"P(x)=(x-1)(x-(1-i))(x-(1+i))"

"P(x)=(x-1)(x^2-2x+2)"

"P(x)=x^3-2x^2+2x-x^2+2x-2"

"P(x)=x^3-3x^2+4x-2"